estimated number of deer in the park.Jared wants to estimate the number of deer in a 600 acre state park. When driving along the one mile side of the road that is on one side of the park, he can see an average of 50 yards into the park and he counts 7 deer. An acre is 4840 square yards, and a mile is 1760 yards. If the deer are evenly distributed throughout the park, what is the best estimate of the number of deer in the park?
Answer by checkley77(12844)

4840*600=2,904,000 yds^2 is the total area.
1760*50=88,000 is the area surveyed with the 7 deer.
7/88,000=x/2,904,000
88,000x=7*2,904,000
88,000x=20,328,000
x=20,328,000/88,000
x=231 is the estimated number of deer in the park.